will you show me step by step please
2(12-8x)=x-11x

Answer by rfer(16322)

2(12-8x)=x-11x
() first
24-16x=x-11x
combine like terms on right
24-16x=-10x
add 16x to both sides
24-16x+16x=-10x+16x
24=6x
divide by 6
x=4